OpenStudy (anonymous):

simplify 6. (x + 6)(x - 6)

OpenStudy (dape):

\[ (a+b)(a-b) = a^2-b^2 \] gives \[ (x+6)(x-6) = x^2-6^2 = x^2-36 \]

OpenStudy (anonymous):

x^2 - 6^2 x^2 - 36

OpenStudy (anonymous):

7. (2x - 7)2

OpenStudy (dape):

\[ (2x-7)*2 = 2x*2-7*2 = 4x-14 \]

OpenStudy (anonymous):

(2x)^2 - 2(2x)(7) + (7)^2 4x^2 -28x + 49

OpenStudy (anonymous):

u use (a - b)^2 = a^2 - 2ab + b^2

OpenStudy (anonymous):

8. (2x + 5)(3x - 7)